OpenStudy (anonymous):

I understand how to do this when the x is on top but stuck on how to do it with it on the bottom

OpenStudy (anonymous):

If \[f(x)=\int\limits_{x}^{17}t^2dt\] then f'(x) =

OpenStudy (anonymous):

t^2

OpenStudy (anonymous):

oh wait maybe not.

OpenStudy (anonymous):

can't you just put in a negative and switch the lims of int.?

OpenStudy (anonymous):

\[ f(x)=\int\limits_{x}^{17}t^2dt = -\int\limits_{17}^{x}t^2dt \]

OpenStudy (anonymous):

duh, ok so then how do I use the 17? the last one was 0 on the bottom and didn't matter

OpenStudy (anonymous):

You could also consider that: \[ \int\limits_{x}^{17}t^2dt = \int\limits_{0}^{17}t^2dt - \int\limits_{0}^{x}t^2dt \]Since: \[ \frac{d}{dx}\int\limits_{0}^{17}t^2dt = 0 \]It would still work out.

OpenStudy (anonymous):

do it like you would a normal integration problem (-t^3/3) from (17 to x)

OpenStudy (anonymous):

then differentiate

OpenStudy (anonymous):

@heradog Since \(17\) is constant with respect to \(x\), it doesn't matter.

OpenStudy (anonymous):

The bottom can be any constant with respect to \(x\) and it still works.

OpenStudy (anonymous):

I didn't think it did, but where does -t^3/3 come from?

OpenStudy (anonymous):

It's an anti-derivative of \(t^2\), don't worry about it because you don't need to use it here.

OpenStudy (anonymous):

Just use \[ g(x) = \int_a^x f(t)dt \implies g'(x) = f(x) \]

OpenStudy (anonymous):

I got the right answer of -x^2 which is where I was hung up. I thought x^2 was the answer but because it was switched it had to be a negative. I've got this one now, so can you help me with one that's more complicated?